# Knowledge Entropy Decay during Language Model Pretraining Hinders New Knowledge Acquisition

### v1

@@ 4.2 Knowledge Acquisition and Retention Decreases across Pretraining Stage
Figure 4 shows the performance of OLMo 1B and 7B models  [^10] For this experiment, we used PubMed as a training corpus, but we also experiment with C4 as a training corpus, which presumably exhibits more similar distribution to pretraining corpus, Dolma. The comparison of the two corpus and the results can be found in Appendix B.1 and B.4.2 corpus from various stages of pretraining as an initial state. We observe that models in the final stage of pretraining struggle more with acquiring new knowledge $\mathcal{A}(\theta)$ and exhibit greater forgetting $\mathcal{F}(\theta)$. As shown in Figure 4, continually training the models at the mid-point of the pretraining as the initial checkpoint tends to yield the best performance in knowledge probing and downstream tasks compared to both models from the initial and final stage of pretraining. While early-stage models demonstrate high knowledge acquisition with minimal forgetting, their overall performance is limited by weaker language modeling capabilities. Conversely, later-stage models, despite being trained on larger datasets, exhibit lower knowledge acquisition and higher rates of forgetting, resulting in lower overall performance compared to the mid-stage models. This aligns with previous research suggesting that a model in the final stage of pretraining tends to struggle when learning new knowledge, showing a tradeoff between plasticity and stability (Dohare et al., 2024; Biesialska et al., 2020; Jang et al., 2022). Therefore, we suggest that using a mid-stage checkpoint strikes a good balance, making it a practical choice as an initial starting point for further training to inject new knowledge.
@@ 4.3 Resuscitating inactive memory vectors increases knowledge acquisition
We observe a strong correlation [^11] The Pearson correlation between knowledge entropy and knowledge acquisition is 0.94, and with forgetting, it is -0.96. Both correlations are statistically significant, with p-values of 6e-5 and 1e-5, respectively. between the trend of knowledge entropy (Figure 3) and the model’s ability to acquire and retain knowledge (Figure 4). We assume that the model’s increasing reliance on a limited set of memory vectors (decrease in knowledge entropy) leads to more frequent updates to these vectors, making it difficult to acquire new knowledge and resulting in a higher rate of forgetting. To test this assumption, we conduct experiments where we artificially increase the activity or resuscitate previously inactive memory vectors.
@@
To resuscitate inactive memory vectors, we modify the up-projection matrix $K$ which engages with producing memory coefficients $\bar{C}$ (notations from Equation 2). Specifically, as shown in Algorithm 1, we identify the lowest $p\%$ (resuscitation ratio) of memory coefficients and apply a multiplier $m$ to parameters in $K$ that are associated with these $p\%$. Multiplier $m$ can be any numbers; in this experiment, we divide the mean coefficient value of each layer by the respective coefficient value $\bar{c}_{i}^{l}$ at each identified position $i$ at layer $l$, and then multiply the result by an amplifying factor $q$. By varying the value of $q$, we control the degree of resuscitation applied to the $p\%$ low-activation coefficients, thereby influencing the magnitude of the average coefficient and corresponding size of the parameter updates.
Figure 5 shows the knowledge acquisition and forgetting rates and Figure 5 presents the knowledge probe and downstream task performance after continual learning with various resuscitation configurations. For the experiment, we fix $p$ to 50 with varying $q$ and use the OLMo checkpoint at the last step of pretraining. Results show that when $q$ is set to 1 or greater, it generally yields better performance in both knowledge acquisition and retention compared to the original model. In contrast, when $q$ is set to 0.5, which further reduces already inactive memory coefficients, both acquisition and retention declines suggesting that concentrating parameter updates more heavily on already active locations led to sparser updates, ultimately impairing overall performance. These results suggest that having a narrower active memory vector (low knowledge entropy) tends to reduce the model’s capacity to acquire new knowledge and increases knowledge forgetting.
@@
2:Scaled up-projection matrix $K$ using computed multiplier $m$
@@
7:     Compute the scaling factor multiplier $m$ for coefficients in layer $l$:
$m=\frac{\text{mean}(C)}{C}\times q$
@@
$K_{l}[\text{idx},:]\hskip 5.0pt\times=\hskip 3.00003ptm[\text{idx}]$
@@
Further experiments with fixed $q$ and varying $p$ shows that increasing $p$ to activate a larger portion of inactive parameters generally led to improved performance. Detailed result of this configuration is in Appendix B.5. We also analyze how the result changes when using models from different stages of pretraining as the original model. We could see that the trend persists over different checkpoints from the final-stage of pretraining (554k). However, the effect of the resuscitation becomes more pronounced as the original model progresses to later stages of pretraining. Detailed results are in Appendix B.6.
Our result indicates that resuscitating inactive memory vectors of final-stage models tends to enhance knowledge acquisition and overall performance than the final-stage model alone. However, we observe that the performance tends to be lower compared to models from the pretraining step with similar knowledge entropy, such as the mid-stage model. This suggests that applying linear scaling to a subset of specific layers alone is insufficient to induce fundamental behavioral changes in the model. In other words, to restore a model that has lost its plasticity (Dohare et al., 2024) to its previous state, more fundamental and alternative approaches are required. Further exploring methods for effectively modifying the parameters would be an interesting direction for future work.
@@ 5 Conclusion
In this work, we examine how large language models’ ability to broadly integrate their parametric knowledge (measured by knowledge entropy) changes throughout pretraining and how these changes affect knowledge acquisition and forgetting in a continual learning setup. Our findings reveal a strong correlation between knowledge entropy and the model’s capacity to acquire and retain knowledge. Models in the final stages of pretraining tend to exhibit narrower integration of memory vectors, leading to lower knowledge entropy, which negatively impacts both knowledge acquisition and retention. Interestingly, we could see that artificially increasing knowledge entropy by modifying the parameters of final-stage models tends to improve these capabilities. Based on our analysis, we suggest that models from the mid-stage of pretraining offer a good balance between knowledge acquisition, retention, and overall performance, making them a good choice for further training to introduce new knowledge.
@@ 6 Limitation & Future Work
Due to computational constraints, our study measures knowledge acquisition and forgetting in a continual learning setup. Future work could explore whether these behaviors also occur during the pretraining phase. We focused on OLMo 1B and 7B models, as they are the only models that publicly provide intermediate pretraining checkpoints and demonstrate strong performance (Sun & Dredze, 2024; Chang et al., 2024). Extending this investigation to other models would be a valuable direction for further research. Our resuscitation method, which arbitrarily modifies model parameters to test our hypothesis, showed promising results in improving knowledge acquisition and retention. However, performance tended to decline when resuscitating models in their initial or mid-stages. This suggests that more refined methods for resuscitating model parameters—ones that avoid random modification and preserve language modeling capabilities—could yield better outcomes. Additionally, while we observed that models in the mid-stage of pretraining strike a good balance for further training on tasks that involve acquiring new knowledge, defining the mid-point precisely remains an open question. In this study, we approximated the mid-point as 50% of the learning rate schedule.
@@ Appendix A Knowledge Entropy
A.1 knowledge entropy
@@
Figure 8 shows how knowledge entropy changes during pretraining by layer. Knowledge entropy consistently decreases in every layer, with the most significant reduction occuring in the last layer, which closely resembles the output distribution right before the token prediction. OLMo-7B also shows similar trend with 1B model.
A.2 Entropy of Attention Layers
Inspired by previous research that emphasizes attention layer’s role of attribute extraction (Geva et al., 2023), we also measure attention entropy $\mathcal{H}(\theta_{\text{att}})$ similarly to Kumar & Sarawagi (2019) [^12] Kumar & Sarawagi (2019) measures entropy using attention weights from cross-attention in an encoder-decoder architecture. However, as we employ a decoder-only model, we modify the equation to use attention weights from self-attention.. Attention weights, output of softmax normalization after key-query-value operation, can be interpreted as weight assigned to the previous tokens. As attention weight for each token position in each attention head is probability distribution(summing to 1), calculating entropy follows normal entropy formula. Then, layer-wise entropy $\mathcal{H}(\theta_{\text{att}}^{l})$ is averaged over token position and attention heads and attention entropy $\mathcal{H}(\theta_{\text{att}})$ is the sum of layer-wise entropy $\mathcal{H}(\theta_{\text{att}}^{l})$. Following the notations from Geva et al. (2023), attention entropy is calculated as:
@@ A.2 Entropy of Attention Layers
where $\mathbf{A}^{h,l,n}\in\mathbb{R}^{T_{n}\times T_{n}}$ represents the attention weights of the $h$-th attention head in layer $l$ for $n$-th instance, $T_{n}$ is the sequence length of the $n$-th instance in the training dataset $\mathcal{D}$, N denotes the number of attention heads, and $L$ denotes the number of layers in the model.
A.3 Entropy of Next Token Prediction
The entropy of next token prediction (Vazhentsev et al., ; Geng et al., 2024; Malinin & Gales, 2021) is defined as $\mathcal{H}(\theta_{\text{ntp}}^{n,j})=-\sum_{i=1}^{|\textit{V}|}p_{i}\cdot\log(p_{i})$, where $p_{i}$ represents the probability of the $i$-th token. This is then averaged over the sequence length ($T_{n}$) and the dataset size ($|\mathcal{D}|$).
@@ B.1 Datasets
In this section, we share a brief description over the datasets we used. For continual knowledge learning, we experiment over PubMed and C4. The PubMed dataset consists of biomedical literature abstracts from the PubMed database, containing articles across a wide range of topics in medicine and biology. The C4 (Colossal Clean Crawled Corpus) dataset (Raffel et al., 2020) is a large-scale, preprocessed collection of text scraped from the web, designed to be a clean and diverse representation of natural language.
@@
To measure knowledge acquisition of language model, we use the fictional knowledge dataset (Chang et al., 2024), which is designed to assess how well LLMs acquire factual knowledge during pretraining. This dataset includes 130 paragraphs  [^13] While Chang et al. (2024) utilizes 120 paragraphs, comprised of 40 for paraphrase setup, 40 for duplicate, and 40 for once, we utilized 130 paragraphs, the whole original data. We scaled up the number of paragraphs with paraphrases to 70, utilizing GPT-4 following Chang et al. (2024) with fictional yet realistic entities (injected knowledge) and 1,950 probes, with three different acquisition depths across five different questions for each knowledge instance, resulting in a total of 15 different probes for each sequence in the corpus. The injected knowledge is incorporated into the training corpus during continual learning, with updates occurring every 160 steps.
@@
To measure forgetting rate, we evaluate over 6 downstream datasets.
  • SciQ: multiple-choice question-answering dataset consisting of over 13,000 science exam-style questions, covering subjects like physics, chemistry, biology, and earth science

### v3

@@ 4.2 Knowledge Acquisition and Retention Decreases across Pretraining Stage
Figure 4 shows the performance of OLMo 1B and 7B models  [^10] For this experiment, we used PubMed as a training corpus, but we also experiment with C4 as a training corpus, which presumably exhibits more similar distribution to the pretraining corpus, Dolma. The comparison of the two corpora and the results can be found in Appendix B.1 and B.4.2 from various stages of pretraining as an initial state. We observe that models in the final stages of pretraining struggle more with acquiring new knowledge $\mathcal{A}(\theta)$ and exhibit greater forgetting $\mathcal{F}(\theta)$. As shown in Figure 4, continually training the models at the mid-point of the pretraining as the initial checkpoint tends to yield the best performance in knowledge probing and downstream tasks compared to both models from the initial and final stages of pretraining. While early-stage models demonstrate high knowledge acquisition with minimal forgetting, their overall performance is limited by weaker language modeling capabilities. Conversely, later-stage models, despite being trained on larger datasets, exhibit lower knowledge acquisition and higher rates of forgetting, resulting in lower overall performance compared to the mid-stage models. This aligns with previous research suggesting that a model in the final stage of pretraining tends to struggle when learning new knowledge, showing a trade-off between plasticity and stability (Dohare et al., 2024; Biesialska et al., 2020; Jang et al., 2022). Therefore, we suggest that using a mid-stage checkpoint strikes a good balance, making it a practical choice as an initial starting point for further training to inject new knowledge.
@@ 4.3 Resuscitating Inactive Memory Vectors Increases Knowledge Acquisition
We observe a strong correlation [^11] The Pearson correlation between knowledge entropy and knowledge acquisition is 0.94, and with forgetting, it is -0.96. Both correlations are statistically significant, with p-values of 6e-5 and 1e-5, respectively. between the trend of knowledge entropy (Figure 3) and the model’s ability to acquire and retain knowledge (Figure 4). We assume that the model’s increasing reliance on a limited set of memory vectors (as indicated by a decrease in knowledge entropy) leads to more frequent updates to these vectors, making it difficult to acquire new knowledge and resulting in a higher rate of forgetting. (The intuition behind this hypothesis can be found in Appendix A.1) To test this assumption, we conduct experiments where we artificially increase the activity or resuscitate previously inactive memory vectors.
@@
To resuscitate inactive memory vectors, we modify the up-projection matrix $K$ which engages with producing memory coefficients $\bar{C}$ (notations from Equation 2). Specifically, as shown in Algorithm 1, we identify the lowest $p\%$ (resuscitation ratio) of memory coefficients and apply a multiplier $u$ to parameters in $K$ that are associated with these $p\%$. Multiplier $u$ can take any value; in this experiment, we divide the mean coefficient value of each layer by the respective coefficient value $\bar{c}_{i}^{l}$ at each identified position $i$ at layer $l$, and then multiply the result by an amplifying factor $q$. By varying the value of $q$, we control the degree of resuscitation applied to the $p\%$ low-activation coefficients, thereby influencing the magnitude of the average coefficient and the corresponding size of the parameter updates.
Figure 5 shows the knowledge acquisition and forgetting rates and Figure 5 presents the knowledge probe and downstream task performance after continual learning with various resuscitation configurations. For the experiment, we fix $p$ to 50 with varying $q$ and use the OLMo checkpoint at the last step of pretraining. Results show that when $q$ is set to 1 or greater, it generally yields better performance in both knowledge acquisition and retention compared to the original model. In contrast, when $q$ is set to 0.5, which further reduces already inactive memory coefficients, both acquisition and retention decline suggesting that concentrating parameter updates more heavily on already active locations leads to sparser updates, ultimately impairing overall performance. These results suggest that having a narrower active memory vector (low knowledge entropy) tends to reduce the model’s capacity to acquire new knowledge and increases knowledge forgetting.
Further experiments with fixed $q$ and varying $p$ show that increasing $p$ to activate a larger portion of inactive parameters generally leads to improved performance. Detailed result of this configuration is in Appendix B.6. We also analyze how the result changes when using models from different stages of pretraining as the original model. The trend holds consistently across different checkpoints of pretraining. However, the effect of the resuscitation becomes more pronounced as the original model progresses to later stages of pretraining. Detailed results are included in Appendix B.7.
Our result indicates that resuscitating inactive memory vectors of final-stage models tends to enhance knowledge acquisition and overall performance compared to the unmodified final-stage model. However, we observe that the performance remains lower compared to models from the pretraining step with similar knowledge entropy, such as the mid-stage model. This suggests that applying linear scaling to a subset of specific layers alone is insufficient to induce fundamental behavioral changes in the model. In other words, to restore a model that has lost its plasticity (Dohare et al., 2024) to its previous state, more fundamental and alternative approaches are required. Further exploring methods for effectively modifying the parameters would be an interesting direction for future work.
@@
2:Scaled up-projection matrix $K$ using computed multiplier $u$
@@
7:     Compute the multiplier $u$ for coefficients in layer $l$:
$u=\frac{\text{mean}(C)}{C}\times q$
@@
$K_{l}[\text{idx},:]\hskip 5.0pt\times=\hskip 3.00003ptu[\text{idx}]$
@@ 5 Conclusion
In this work, we examine how large language models’ ability to broadly integrate their parametric knowledge (measured by knowledge entropy) changes throughout pretraining and how these changes affect knowledge acquisition and forgetting in a continual learning setup. Our findings reveal a strong correlation between knowledge entropy and the model’s capacity to acquire and retain knowledge. Models in the final stages of pretraining tend to exhibit narrower integration of memory vectors, leading to lower knowledge entropy, which negatively impacts both knowledge acquisition and retention. Interestingly, artificially increasing knowledge entropy by modifying the parameters of final-stage models tends to improve these capabilities. Based on our analysis, we suggest that models from the mid-stage of pretraining offer a good balance between knowledge acquisition, retention, and overall performance, making them a good choice for further training to introduce new knowledge.
@@ 6 Limitation & Future Work
Due to computational constraints, our study measures knowledge acquisition and forgetting in a continual learning setup. Future work could explore whether these behaviors also manifest during the pretraining phase. We focused on OLMo 1B and 7B models, as they are the only models that publicly provide intermediate pretraining checkpoints and demonstrate strong performance (Sun & Dredze, 2024; Chang et al., 2024). Extending this investigation to other models would be a valuable direction for further research. Our resuscitation method, which arbitrarily modifies model parameters to test our hypothesis, showed promising results in improving knowledge acquisition and retention. However, performance tended to decline when resuscitating models at their initial or mid-stages. This suggests that more refined methods for resuscitating model parameters—ones that avoid random modification and preserve language modeling capabilities—could yield better outcomes. Additionally, while we observed that models in the mid-stage of pretraining strike a good balance for further training on tasks that involve acquiring new knowledge, defining the mid-point precisely remains an open question. In this study, we approximated the mid-point as 50% of the learning rate schedule.
@@ Acknowledgments
This work was supported by LG AI Research grant (Self-improving logical reasoning capabilities of LLMs, 2024, 50%) and the Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2024-00397966, Development of a Cybersecurity Specialized RAG-based sLLM Model for Suppressing Gen-AI Malfunctions and Construction of a Publicly Demonstration Platform, 20%; No.2022-0-00113, Developing a Sustainable Collaborative Multi-modal Lifelong Learning Framework, 20%; No.RS-2021-II212068, Artificial Intelligence Innovation Hub, 10%)
@@ Appendix A Knowledge Entropy
A.1 Intuition Behind the Definition of Knowledge Entropy
The mechanism behind the relationship between decreasing knowledge entropy and the ability to acquire and retain knowledge during pretraining is that the coefficients in the linear combination of memory vectors determine how the corresponding memory vectors are updated. As defined in Equation 1, the output of the feed-forward layers (FFNs) is a linear combination of memory vectors $\mathbf{v}_{i=1,\cdots,m}\in\mathbb{R}^{d}$, the row vectors of $V\in\mathbb{R}^{m\times d}$, where the coefficients $c_{i}$ are given by $f(\mathbf{x}\cdot K^{T})$. In other words, $FFN(x)=c_{1}\mathbf{v_{1}}+c_{2}\mathbf{v_{2}}+\cdots+c_{m}\mathbf{v_{m}}$. Within a given layer, since the operations beyond the FFNs and the input to the FFNs remain consistent across all memory vectors, the coefficients act as scaling factors for the gradient by the chain rule. During training, the gradient $\frac{\partial L}{\partial\mathbf{v_{i,j}}}$ for $i=1,2,\ldots,m$ and $j=1,2,\ldots,d$ can be decomposed as:
$\frac{\partial L}{\partial\mathbf{v_{i,j}}}=\frac{\partial L}{\partial FFN(x)}\cdot\frac{\partial FFN(x)}{\partial v_{i,j}}$
Here, $\frac{\partial L}{\partial FFN(x)}$ is the same for all $\mathbf{v_{i}}$, meaning that the relative magnitude of the gradient depends on $\frac{\partial FFN(x)}{\partial v_{i,j}}=c_{i}$. Thus, larger coefficients result in proportionally larger gradients being applied to the corresponding memory vectors, amplifying their updates during backpropagation.
As pretraining progresses, a spikier coefficient distribution—captured by decreasing knowledge entropy—implies that gradient updates become increasingly concentrated on specific positions where the average coefficients are larger. This centralization can affect the model’s ability to evenly utilize its memory capacity, thereby impacting knowledge acquisition and retention.
A.2 Knowledge Entropy
Figure 6: Knowledge Entropy (y-axis) measured with Dolma on different pretraining stages of Pythia 1.4B model.
@@ A.2 Knowledge Entropy
Does the choice of model change the trend?
To assess the generalizability of the trend observed in Figure 3, we conducted experiments on the knowledge entropy trend using Pythia 1.4B model (Biderman et al., 2023). As shown in Figure 7, knowledge entropy measured with the Pythia model also tends to decrease as pretraining progresses.
@@
Figure 8 shows how knowledge entropy changes by layer during pretraining. Knowledge entropy consistently decreases in every layer, with the most significant reduction occurring in the last layer, which closely resembles the output distribution right before the token prediction. OLMo-7B model also shows a similar trend to 1B model.
A.3 Entropy of Attention Layers
Inspired by previous research that emphasizes the attention layer’s role of attribute extraction (Geva et al., 2023), we also measure attention entropy $\mathcal{H}(\theta_{\text{att}})$ similarly to Kumar & Sarawagi (2019) [^12] Kumar & Sarawagi (2019) measures entropy using attention weights from cross-attention in an encoder-decoder architecture. However, as we employ a decoder-only model, we modify the equation to use attention weights from self-attention.. Attention weights, which are the output of softmax normalization after the key-query-value operation, can be interpreted as weight assigned to the previous tokens. As the attention weight for each token position in each attention head forms a probability distribution(summing to 1), calculating entropy follows the normal entropy formula. Then, layer-wise entropy $\mathcal{H}(\theta_{\text{att}}^{l})$ is averaged over token position and attention heads and attention entropy $\mathcal{H}(\theta_{\text{att}})$ is the sum of layer-wise entropy $\mathcal{H}(\theta_{\text{att}}^{l})$. Following the notations from Geva et al. (2023), attention entropy is calculated as:
@@ A.3 Entropy of Attention Layers
where $\mathbf{A}^{h,l,n}\in\mathbb{R}^{T_{n}\times T_{n}}$ represents the attention weights of the $h$-th attention head in layer $l$ for the $n$-th instance, $T_{n}$ is the sequence length of the $n$-th instance in the training dataset $\mathcal{D}$, N denotes the number of attention heads, and $L$ denotes the number of layers in the model.
A.4 Entropy of Next Token Prediction
The entropy of next token prediction (Vazhentsev et al., ; Geng et al., 2024; Malinin & Gales, 2021) is defined as $\mathcal{H}(\theta_{\text{ntp}}^{n,j})=-\sum_{i=1}^{|\textit{V}|}p_{i}\cdot\log(p_{i})$, where $p_{i}$ represents the probability of the $i$-th token. This value is then averaged over the sequence length ($T_{n}$) and the dataset size ($|\mathcal{D}|$).
@@ B.1 Datasets
In this section, we share a brief description of the datasets we used. For continual knowledge learning, we experiment with PubMed and C4. The PubMed dataset consists of biomedical literature abstracts from the PubMed database, containing articles across a wide range of topics in medicine and biology. The C4 (Colossal Clean Crawled Corpus) dataset (Raffel et al., 2020) is a large-scale, preprocessed collection of text scraped from the web, designed to be a clean and diverse representation of natural language.
@@
To measure knowledge acquisition of language model, we use the fictional knowledge dataset (Chang et al., 2024), which is designed to assess how well LLMs acquire factual knowledge during pretraining. This dataset includes 130 paragraphs  [^13] While Chang et al. (2024) utilizes 120 paragraphs, comprising 40 for paraphrase setup, 40 for duplicate, and 40 for once, we utilized 130 paragraphs, the whole original data. We scaled up the number of paragraphs with paraphrases to 70, utilizing GPT-4 following Chang et al. (2024)  presented in a Wikipedia-style format with fictional yet realistic entities (injected knowledge) and 1,950 probes which are cloze-task-style sentences to query the information within the corpus. The final span of each probe, referred to as the target span, is used to evaluate the model’s prediction probability, which serves as a measure of knowledge acquisition performance.
In Chang et al. (2024), the probes are divided into three levels of difficulty, with five sentences created for each level. This results in 15 probes per corpus. The difficulty levels are as follows: 1) Memorization probes directly ask about sentences explicitly present in the fictional corpus. 2) Semantic generalization probes are paraphrased versions of the memorization probes to test the model’s understanding of meaning beyond surface forms. 3) Compositional generalization probes are designed to assess whether the model can integrate multiple pieces of knowledge from the fictional corpus. The injected knowledge is incorporated into the training corpus during continual learning, with updates occurring every 160 steps.
@@
To measure the forgetting rate, we evaluate on 6 downstream datasets.
  • SciQ: multiple-choice question-answering dataset consisting of over 13,000 science exam-style questions, covering subjects such as physics, chemistry, biology, and earth science